OpenStudy (anonymous):

a) The average height of sunflowers in a field is 64 inches with a standard deviation of 3.5 inches. Describe a normal curve for the distribution, including the values on the horizontal axis at one, two, and three standard deviations from the mean. b) If there are 3,000 plants in the field, approximately how many will be taller than 71 inches?

OpenStudy (anonymous):

were you given any options?

OpenStudy (anonymous):

no. I have to show my work in a blank thingy.

OpenStudy (anonymous):

fun. ok... hold on... let me try to do it first.

OpenStudy (anonymous):

x = 64 y = 3.5 a) x - 3y = 53.5 x - 2y = 57.0 x - y= 60.5 x = 64.0 x + y = 67.5 x + 2y = 71.0 x + 3y = 74.5 ? i think that is how....

OpenStudy (anonymous):

and b is (71 - 64)/3.5 = 2.0 z = 0.9772 n = 3000*(1 - 0.9772) ≈ 68

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

yep. you can private message me if you need more help... it may be easier. I check that more.

OpenStudy (anonymous):

Okay. One question, why did you use 71?

OpenStudy (anonymous):

because it says 71 inches

OpenStudy (anonymous):

Oh. oops. I forgot about that. Okay. Tysm!

OpenStudy (anonymous):

anytime.

OpenStudy (anonymous):

One more i'm stuck on. John puts the letters of the word SCIENCE on pieces of paper and puts the pieces in a bag. He draws two letters without looking. The first piece of paper is not replaced before the second piece is drawn. What is the probability that John draws a C twice? A. 2/49 b. 1/7 c. 1/14 d. 1/21

OpenStudy (anonymous):

@jtbulin

OpenStudy (anonymous):

ok.. this is not bad.. so there is 2 c's in sCienCe. There is 7 letters.

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

to do this, it is kind of sucky though. it is 7P2. that means that P is a permutation, if you plus it into a calculator, in this case, just use http://www.wolframalpha.com/input/?i=7P2 then your answer is 49 choices total. that means that it would have to have that in the denominator. It is A, I am pretty sure. I did the same problem last year...

OpenStudy (anonymous):

Oh okay. I was thinking that because 2 would be the numerator, but I wasn't sure were the 49 came from. But thanks so much... again.

OpenStudy (anonymous):

no problem